# Generative Regression Based Watch Time Prediction for Video Recommendation: Model and Performance

### v1

@@ Abstract.
Watch time prediction (WTP) has emerged as a pivotal task in short video recommendation systems, designed to encapsulate user interests. Predicting users’ watch times on videos often encounters challenges, including wide value ranges and imbalanced data distributions, which can lead to significant bias when directly regressing watch time. Recent studies have tried to tackle these issues by converting the continuous watch time estimation into an ordinal classification task. While these methods are somewhat effective, they exhibit notable limitations. For instance, the discretization process often hinges heavily on bucket partitioning, which inherently compromises prediction accuracy. Moreover, the relationships existing between various partition intervals are seldom fully exploited, thereby ignoring the potential for effective error correction.
Inspired by language modeling, we propose a novel Generative Regression (GR) paradigm for WTP based on sequence generation. This approach employs structural discretization to enable the lossless reconstruction of original values while maintaining prediction fidelity. By formulating the prediction problem as a numerical-to-sequence mapping, and with meticulously designed vocabulary and label encodings, each watch time is transformed into a sequence of tokens. To expedite model training, we introduce the curriculum learning with an embedding mixup strategy which can mitigate training-and-inference inconsistency associated with teacher forcing. We evaluate our method against state-of-the-art approaches on four public datasets and one industrial dataset. We also perform online A/B testing on Kuaishou, a leading video app with about 400 million DAUs, to demonstrate the real-world efficacy of our method. The results conclusively show that GR outperforms existing techniques significantly. Furthermore, we successfully apply GR to another regression task in recommendation systems, i.e., Lifetime Value (LTV) prediction, which highlights its potential as a novel and effective solution to general regression challenges.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The Web Conference 2025; April 28–May 2, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems

### v2

# Sequence Generation Modeling for Continuous Value Prediction

@@ Abstract.
Continuous value prediction (CVP) plays a crucial role in short video recommendation, capturing user preferences through precise numerical estimations. However, traditional regression-based methods often struggle with challenges like wide value ranges and imbalanced data, leading to prediction bias. While ordinal classification approaches have been introduced to address these issues, their reliance on discretization reduces accuracy and overlooks inherent relationships between intervals. To overcome these limitations, we introduce a novel Generative Regression (GR) framework for CVP, inspired by sequence generation techniques in language modeling. Our method transforms numerical values into token sequences through structural discretization, preserving original data fidelity while improving prediction precision. Leveraging a carefully crafted vocabulary and label encoding, GR employs curriculum learning with an embedding mixup strategy to bridge training-inference gaps. Experimental evaluations on four public datasets and one large-scale industrial dataset validate the superiority of GR over existing methods. Real-world A/B tests on Kuaishou, a leading video platform, further demonstrate its practical effectiveness. Additionally, GR proves adaptable to other regression tasks, such as Lifetime Value (LTV) prediction, showcasing its potential as a robust solution for diverse CVP challenges.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems